OpenStudy (anonymous):

Solve (7x + 2)^2 = 64

OpenStudy (jhannybean):

What do you get when you take the square root of both sides?

OpenStudy (jhannybean):

\[\sqrt{(7x+2)^2}= \pm \sqrt{64}\]

OpenStudy (jhannybean):

So when you take the square root of a square, you get the function inside, this would simply be 7x+2.

OpenStudy (jhannybean):

I'll leave the rest to you :)